# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v3

@@ 6.1 Rare Events and the Value Function
We are concerned with rare event states if they have a sizeable impact on the user experienced costs, i.e. if these states collectively have an impact on the value function. The value function $V^{\pi}(s):S\rightarrow\mathbb{R}$ for the policy $\pi$ is the expected-return of state $s$, i.e. it indicates how “good” (cost wise) a state is (based on the discounted sum of future rewards), when $\pi$ is used. The discount factor $\gamma$ indicates how much the network operator cares about future vs current rewards.
@@
Likewise, defining $U^{\pi}(s)$ as the collective contribution of the non-rare states (states in $S\backslash T$) towards the value of state $s$, given a fixed policy $\pi$, we have :
@@
$=\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\sum_{a^{\prime}\in A}\pi(a^{\prime}|s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})]$
@@ 6.2 FIRE-ImRE: Q-Learning in Light of Rare Events
In this subsection, we present our importance sampling based failure-adaptive algorithm. We use a digital twin-based simulator in order to avoid the large failure costs experienced during server failures in online scenarios, and the converged policy trained in the digital twin is then applied to online scenarios, in which rare events happen at their natural probabilities. Our algorithm performs importance sampling while increasing the rare event probabilities from $\epsilon(s)$ to $\hat{\epsilon}(s)$. The transition probabilities $h(\tilde{s}|s,a)$, which mostly depend on the user mobility probabilities (Eq. (3)), are obtained from historical data. We use importance sampling because if rare events such as server failures are sampled at their natural probabilities, the reinforcement learning algorithm is not able to converge to the optimal policy; in particular, it may decide that the storage cost isn’t worth having backups. This can be catastrophic in the event of failures. In our numerical simulations in Section 9, we show that unlike our algorithm, a traditional RL algorithm is unable to train an optimal policy which avoids experiencing high cost in the event of failures (Figs. 4 - 6).
@@
Updating at the state action $(s,a)$ level may lead to slower convergence, especially with an increasing number of access points. For example, with 9 access points, there will be 360 states and 90 actions, leading to 32400 state-action combinations. We can reduce this complexity by exploiting our earlier observation that the probability of rare events (i.e., server failures) happening is independent of the actions chosen. Therefore, we propose updating $T$ and $U$ (the contributions of the rare events and non-rare events respectively towards the Q-value) at the state level:
@@
$\hat{\epsilon}(s)\leftarrow min(max(\delta,\frac{|T(s)|}{|T(s)|+|U(s)|}),1-\delta).$ (30)
@@ 6.3 Boundedness and Convergence Properties
The main idea of the proof is as follows: We rewrite Algorithm 1’s update equation, through adding and subtracting terms. We rewrite the equation in the form of $\hat{Q}^{t+1}(s,a)=\hat{Q}^{t}(s,a)+\alpha^{t}[F_{s,a}(\hat{Q}^{t})-\hat{Q}^{t}(s^{t},a^{t})+M_{s}(t)]$. Invoking Theorem 1 of [33] on stochastic approximation algorithms, we finally prove convergence by showing that the conditions for the sequence $F_{s,a}(\hat{Q})$ to be bounded are met. We do this by showing that $||M(t)||$ is bounded, and that the property $||F(\hat{Q})||_{\infty}\leq b||\hat{Q}||_{\infty}+d$ holds, where $||\quad||_{\infty}$ is the supremum norm.
@@
$||\mathbb{E}[F_{t}(x)]||\leq\gamma||\triangle_{t}||,\text{where}\ \gamma<1.$
$var[F_{t}(x)]\leq C(1+||\triangle_{t}||^{2}),\ \text{for}\ C>0.$
@@
If the MDP underlying the RL environment is unichain for $\epsilon\in(\delta,1-\delta)$,111$(\delta,1-\delta)$ is the range of values which $\hat{\epsilon}$ takes, as defined in Algorithm 1. If the MDP defined by transition probability $p(s^{\prime}|s,a)$ is unichain for one value in $(\delta,1-\delta)$, it is unichain for all values [14]. for stepsizes satisfying $0\leq\alpha_{t}\leq 1$, $\sum_{t}\alpha_{i}(t)=\infty$ and $\sum_{t}\alpha_{i}^{2}(t)\leq\infty$, and discount factor $\gamma\in(0,1)$, Algorithm 1 converges to the optimal policy $Q^{*}$.
@@
The main idea of the proof is as follows. Firstly, we show that the optimal policy of the true system is a fixed point of the following equation with our simulator’s transition probability $\hat{p}(s^{\prime}|s,a)$ and the importance sampling correction weight $w_{t}(s)$: $Q^{*}(s,a)=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}Q^{*}(s^{\prime},b)].$ We define the operator $\mathbf{H}\hat{Q}=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}(s^{\prime},b)]$, and having shown that that $\mathbf{H}Q^{*}=Q^{*}$, we next show that $\mathbf{H}\hat{Q}$ is a contraction mapping, through showing that $||\mathbf{H}\hat{Q}_{1}-\mathbf{H}\hat{Q}_{2}||_{\infty}\leq\gamma||\hat{Q}_{1}-\hat{Q}_{2}||_{\infty}$. Next, we invoke the result of Corollary 3 on the convergence of random processes. We rewrite the update equation of Algorithm 1, such that it fits the form of the equations in Corollary 3, and show that the conditions of Corollary 3 hold. We do this by letting $G_{t}(s,a)=w^{t}[r(s,a,X(s,a))+\gamma\max_{b}\hat{Q}^{t}(X(s,a),b)]-Q^{*}(s,a)$, and by using the fact that $\mathbf{H}\hat{Q}$ is a contraction mapping, we show that $||\mathbb{E}[G_{t}(s,a)]||_{\infty}\leq\gamma||\hat{Q}^{t}(s,a)-Q^{*}||_{\infty}=\gamma||\triangle_{t}||_{\infty}$, and $\mathbf{var}[G_{t}(s,a)]\leq C(1+||\triangle_{t}||^{2}_{\infty})$.
@@ 7 Scaling to Large State or Action Spaces
Rare Event Adaptive Deep Q-learning algorithm (FIRE-ImDQL, Algorithm 2): We modify and combine the classic deep reinforcement learning algorithm in [37] with our importance sampling based rare events adaptive algorithm (Algorithm 1) by proposing an importance sampling based loss function.
@@
where $\boldsymbol{\theta}^{-}$ are the weights of the target neural network, $\boldsymbol{\theta}$ are the weights of the predicted neural network, which are updated every iteration, and $w_{t}$ is the importance sampling weight. The loss function is optimized via gradient descent in line 13. We are trying to minimize the difference between the target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ and the prediction $Q(s,a,\boldsymbol{\theta})$. The target network is updated only every $K$ iterations (line 16), while the prediction network is updated every iteration. Having the target fixed for awhile helps in having stable training. The key idea which differentiates our algorithm from traditional deep reinforcement learning algorithms is that our target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ is “error corrected” by the importance sampling weight $w_{t}$, and the rare event distribution is updated via $T,U$ and $\hat{\epsilon}$. The variables $T$, $U$, $\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1), through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t+1}$. The actions are also selected in the same manner as Algorithm 1, through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t}$.
@@ 8 Heterogeneous Risk Tolerances
2:$p_{riskAware}\leftarrow\text{softmax}(Q_{riskAware}(s,a))$
3:$p_{riskTaking}\leftarrow\text{softmax}(Q_{riskTaking}(s,a))$
@@
After $T$ time-slots, at risk level $\zeta$, assuming our algorithm is able to prevent the cost of failures with probability $1-f$, it benefited the user to prepare for rare events, if
@@
(c) Convergence of deep Q-learning importance sampling algorithm FIRE-ImDQL.
Figure 3: Convergence of our proposed importance-sampling based rare events adaptive algorithms. Here, FIRE-ImRE and FIRE-ImACRE are applied to the a special case of the scenario (Section 3) where every user has their own service profile (the single user case), and FIRE-ImDQL is applied to the scenario where users share service profiles (Section 4). All three variations of our algorithm converge.
@@ 9 Simulation Experiments
Single User Service Profile Migrations: In our experiments, we have 9 base stations (edge access points). The user is highly mobile across these coverage areas, and its mobility pattern follows a transition matrix. The ns-3 network simulator is used [39] to simulate dynamic channel network conditions, to obtain latency (delay) costs across different user-server location pairs. These latencies are in the range (2,10). As both the latency costs and migration costs are functions of the distance between user and access points, we set the migration cost $m_{ij}=d^{comm}_{i,j}+\varepsilon$, where $\varepsilon\in(-0.5,0.5)$. We set the cost of failures to be -500, representing the high cost in terms of user experience or safety, and the true failure rate as 0.01 [41]. We choose this failure rate, because a higher order of 0.1 will entail that the failures are no longer rare events. And at a lower order of 0.001, the failures would give a smaller expected cost to the system. The learning rate used is $0.05$ and both the actor and critic networks, for both our algorithms and the baselines, have a structure of $[24,48,24]$ nodes per layer.
@@
Multiple User Service Profile Migrations: Here, we consider service profile migration for multiple users, each with their own service profile. We model 13 users mobile across 9 access points. Parameter settings are similar to the setup above. The natural failure rate is set to be 0.01. User mobility patterns are sampled according to a transition probability matrix. We model different server failure types, corresponding to the differing amounts of maintenance resources at each AP, which affect the time taken for the servers to come back online. For APs 1,6, and 8, we let their service downtime be 2 time-slots per failure. For all other APs, we let their service downtime be 1 time-slot per failure. We apply algorithm FIRE-ImDQL, our importance sampling based deep Q learning algorithm in this scenario. In Fig. 5, we see that our FIRE-ImDQL policy outperforms the baselines by achieving a lower average cost during rare states, with a trade-off of a higher average cost in normal (non-rare) states. Overall, our algorithm still achieves a lower cost, when averaged over all timeslots. QL-NIS and the greedy algorithm perform better than the other baselines, as backups are still a potential action for the algorithm in QL-NIS, and the greedy algorithm places one backup at the second most likely location every time-slot. Nevertheless these algorithms do not avoid the rare event cost as much as our algorithm does.
@@
Shared service profiles: In this scenario (Fig. 6), we consider multiple users sharing the same service profile (SP). We model 100 users, with 70 users using SP Type 1, and 30 using SP Type 2. These users are spread over a 2-city map, with 2APs in one city and 3 in another. It has a failure cost of 1000 and failure rate of 1e-2 [41]. If a user’s required access point is in another city, the delay cost between cities will be nearly as large as the failure cost, which is larger than the delay cost when accessing an SP within the same city. We apply FIRE-ImDQL to this scenario and compare its performance with NIS, RES and the greedy baseline. The greedy baseline deploys SPs to the $N$ most probable places w.r.t distribution of user transitions, where we search through all $N$ and plot the results with the lowest costs.

### v4

@@ 6.1 Rare Events and the Value Function
We define rare event states as occasions when service fails: $f^{k}_{ind}(t)=1$ in Section 3 or $f_{i}(t)=1$ in Section 4. We are concerned with rare event states if they have a sizeable impact on the user experienced costs, i.e. if these states collectively have an impact on the value function. The value function $V^{\pi}(s):S\rightarrow\mathbb{R}$ for the policy $\pi$ is the expected-return of state $s$, i.e. it indicates how “good” (cost wise) a state is (based on the discounted sum of future rewards), when $\pi$ is used. The discount factor $\gamma$ indicates how much the network operator cares about future vs current rewards.
@@
[…]
@@
$\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\sum_{a^{\prime}\in A}\pi(a^{\prime}|s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})]$
@@ 6.2 FIRE-ImRE: Q-Learning in Light of Rare Events
In this subsection, we present our importance sampling based failure-adaptive algorithm. We use a digital twin-based simulator in order to avoid the large failure costs experienced during server failures in online scenarios, and the converged policy trained in the digital twin is then applied to online scenarios, in which rare events happen at their natural probabilities. Our algorithm performs importance sampling while increasing the rare event probabilities from $\epsilon(s)$ to $\hat{\epsilon}(s)$. The transition probabilities $h(\tilde{s}|s,a)$, which mostly depend on the user mobility probabilities (Eq. (3)), are obtained from historical data. We use importance sampling because if rare events such as server failures are sampled at their natural probabilities, the RL algorithm is not able to converge to the optimal policy; in particular, it may decide that the storage cost isn’t worth having backups. This can be catastrophic in the event of failures. In our numerical simulations in Section 9, we show that unlike our algorithm, a traditional RL algorithm is unable to train an optimal policy which avoids experiencing high cost in the event of failures (Figs. 5 - 7).
@@
Updating at the state action $(s,a)$ level may lead to slower convergence, especially with an increasing number of APs. For example, with 9 APs, there will be 360 states and 90 actions, leading to 32400 state-action combinations. We can reduce this complexity by exploiting our earlier observation that the probability of rare events (i.e., server failures) happening is independent of the actions chosen. Therefore, we propose updating $T$ and $U$ (the contributions of the rare events and non-rare events respectively towards the Q-value) at the state level:
@@
$\hat{\epsilon}(s)\leftarrow\min(\max(\delta,\frac{|T(s)|}{|T(s)|+|U(s)|}),1-\delta).$ (30)
@@ 6.3 Boundedness and Convergence Properties
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7 Scaling to Large State or Action Spaces
Rare Event Adaptive Deep Q-learning algorithm (FIRE-ImDQL, Algorithm 2): We modify and combine the classic deep RL algorithm in [40] with our importance sampling based rare events adaptive algorithm (Algorithm 1) by proposing an importance sampling based loss function.
@@
where $\boldsymbol{\theta}^{-}$ are the weights of the target neural network, $\boldsymbol{\theta}$ are the weights of the predicted neural network, which are updated every iteration, and $w_{t}$ is the importance sampling weight. The loss function is optimized via gradient descent in line 13. We are trying to minimize the difference between the target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ and the prediction $Q(s,a,\boldsymbol{\theta})$. The target network is updated only every $K$ iterations (line 16), while the prediction network is updated every iteration. Having the target fixed for awhile helps in having stable training. The key idea which differentiates our algorithm from traditional deep RL algorithms is that our target $w_{t}(r^{t+1}+\gamma\max_{a}\hat{Q}(s,a,\boldsymbol{\theta}^{-}))$ is “error corrected” by the importance sampling weight $w_{t}$, and the rare event distribution is updated via $T,U$ and $\hat{\epsilon}$. The variables $T$, $U$, $\hat{\epsilon}$ and $w_{t}$ are updated in the same manner as the tabular algorithm (Algorithm 1), through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t+1}$. The actions are also selected in the same manner as Algorithm 1, through using the output of the prediction net $Q(s,a,\boldsymbol{\theta})$ where the input is state $s^{t}$.
@@
Application to other Settings with Rare Events: Our algorithm can also be applied for other networked and wireless settings, in which there are rare events. An example application is UAV-assisted wireless networks with potential node failures. The failures are: UAV nodes failing (e.g. due to physical damage). The state space could include UAV positions and channel conditions, and the number of users served by each UAV. Possible actions include the user-UAV allocation, the bandwidth allocation, and the decision whether or not to deploying backup UAVs. The environment, state space, action space, and rare events would be specific to the new setting, while our importance-sampling algorithm itself would remain the same.
@@ 8 Heterogeneous Risk Tolerances
Figure 3: Algorithm Framework: Our importance sampling based reinforcement learning algorithms learns offline in a digital twin setup, to avoid experiencing the real cost of rare events. The converged policy is applied to online scenarios where rare events occur at their true rate.
We present our overall algorithm framework in Fig 3. Our importance sampling based reinforcement learning algorithms learns offline in a digital twin setup, to avoid experiencing the real cost of rare events. The converged policy is applied to online scenarios where rare events occur at their true rate. Running the policy once trained is identical to deploying any RL-trained policy. If we use, e.g., tabular RL, this is a table lookup; deep RL may require some inference time through the neural network.
@@
[…]
@@
After $T$ time-slots, at risk level $\zeta$, assuming our algorithm is able to prevent the cost of failures with probability 1f1-f, it benefited the user to prepare for rare events, if
@@
(c) Convergence of deep Q-learning importance sampling algorithm FIRE-ImDQL, shared SP scenario.
(d) Convergence of deep Q-learning algorithm FIRE-ImDQL, individual SP scenario.
Figure 4: Convergence graphs: Here, FIRE-ImRE and FIRE-ImACRE are applied to the a special case of the scenario (Section 3) where every user has their own service profile (the single user case), and FIRE-ImDQL is applied to the scenario where users share service profiles (Section 4). All three variations of our algorithm converge.
@@ 9 Simulation Experiments
Single User Service Profile Migrations: In our experiments, we have 9 base stations (edge APs). The user is highly mobile across these coverage areas, and its mobility pattern follows a transition matrix. We use the same network settings as in [19] to simulate dynamic latencies between user locations and different base stations as users move around the service area: In particular, the ns-3 network simulator is used [42] to simulate dynamic channel network conditions, to obtain these latencies. These latencies are in the range (2,10). As both the latency costs and migration costs are functions of the distance between user and APs, we set the migration cost $m_{ij}=d^{comm}_{i,j}+\varepsilon$, where $\varepsilon\in(-0.5,0.5)$. We set the cost of failures to be -500, representing the high cost in terms of user experience or safety, and the true failure rate as 0.01 (1%) [44]. We choose this failure rate, because if the failure rate is of a higher order (e.g. 15 percent), then we may not need a special failure-aware importance sampling strategy as the ‘failures’ are no longer rare events - they appear often enough for normal RL methods to learn a policy which deals with them. Conversely, if the failure rate is lower (e.g. 0.001=0.1 percent), the failures would give a smaller expected cost on the system (Definition 1 in Section 3.1), so are less important to handle. The learning rate used is $0.05$ and both the actor and critic networks, for both our algorithms and the baselines, have a structure of $[24,48,24]$ nodes per layer.
@@
Multiple User Service Profile Migrations: Here, we consider service profile migration for multiple users, each with their own service profile. We model 13 users mobile across 9 APs. Parameter settings are similar to the setup above. The natural failure rate is set to be 0.01. User mobility patterns are sampled according to a transition probability matrix. We model different server failure types, corresponding to the differing amounts of maintenance resources at each AP, which affect the time taken for the servers to come back online. For APs 1,6, and 8, we let their service downtime be 2 time-slots per failure. For all other APs, we let their service downtime be 1 time-slot per failure. We apply algorithm FIRE-ImDQL, our importance sampling based deep Q learning algorithm in this scenario. In Fig. 6, we see that our FIRE-ImDQL policy outperforms the baselines by achieving a lower average cost during rare states, with a trade-off of a higher average cost in normal (non-rare) states. Overall, our algorithm still achieves a lower cost, when averaged over all timeslots. QL-NIS and the greedy algorithm perform better than the other baselines, as backups are still a potential action for the algorithm in QL-NIS, and the greedy algorithm places one backup at the second most likely location every time-slot. Nevertheless these algorithms do not avoid the rare event cost as much as our algorithm does.
@@
Comparison with PPO: In Table I, we present results in which our algorithm is compared with PPO (proximal policy optimization), for the multiuser scenario, where users have individual service profiles. The results are averaged over 6 runs. It outperforms PPO when the converged policies of both algorithms are compared in an online setting where rare events occur at their true rate: PPO has a higher average cost over time, and higher rare event cost, as there is no importance sampling and hence lack of rare event data to train a policy which is adaptive to rare events.
TABLE I: Comparison of ImDQL with PPO, averaged over 6 runs. For both algorithms, the converged policy is applied in an online scenario where rare events occur at their true rate.
Reward across time, online setting ImDQL PPO
Avg reward across time -18.210 -22.031
Mean RE reward -18.689 -551.79
Mean RE migration cost -6.688 -4.628
Mean RE delay -5.176 -5.342
Mean RE storage cost -5.000 -4.798
Mean RE computing cost -0.9289 -0.889
Mean NS reward -17.279 -14.584
Mean NS migration cost -6.968 -4.656
Mean NS delay -5.311 -5.171
Mean NS storage cost -5.000 -4.757
Mean NS comp cost -0.908 -0.895
Higher failure rates: We have further performed new experiments with higher failure rates, namely 5 and 10% (Tables II and III). Our algorithm is able to handle higher failure rates. Note that if failures are fairly common, then we may not need to use importance sampling to amplify rare events in the training dataset, as the rare failure events would naturally occur at a fairly high rate. Thus, reinforcement learning algorithms without importance sampling should still be able to learn good policies that account for failure costs, as these costs would be present in the training dataset. Indeed, we observe this effect in our results. At a failure rate of 5 percent (Table II), our algorithm still outperforms the baselines, achieving a much lower mean rare events reward. QL-NIS outperforms QL-WBA and QL-RES in the online scenario because backups and rare events are considered in training. At a failure rate of 10 percent (Table III), the performance of QL-NIS improves, as it is easier learn a policy in light of not-as-rare server failures. While they are of the same order, our algorithm still outperforms it.
[…]
Reward across time ImDQL QL-NIS QL-WBA QL-RES
Avg reward across time -19.586 -26.062 -416.322 -578.736
Mean RE reward -20.679 -186.516 -8002.600 -8003.399
Mean RE migration cost -7.663 -6.284 -2.600 -3.399
Mean RE delay -5.189 -4.785 0.000 0.000
Mean RE storage cost -4.999 -4.914 0.000 0.000
Mean RE computing cost -0.957 -1.029 -0.929 -0.993
Mean NS reward -18.394 -15.483 -8.061 -7.804
Mean NS migration cost -8.053 -6.332 -2.609 -3.456
Mean NS delay -5.344 -4.244 -5.452 -4.349
Mean NS storage cost -4.997 -4.907 0.000 0.000
Mean NS computing cost -0.924 -1.038 -0.932 -1.003
[…]
Reward across time ImDQL QL-NIS QL-WBA QL-RES
Avg reward across time -20.218 -18.665 -416.322 -578.736
Mean RE reward -23.414 -42.013 -8002.600 -8003.399
Mean RE migration cost -7.643 -5.568 -2.600 -3.399
Mean RE delay -5.124 -4.652 0.000 0.000
Mean RE storage cost -4.997 -4.984 0.000 0.000
Mean RE computing cost -0.943 -1.027 -0.929 -0.993
Mean NS reward -18.265 -14.698 -8.061 -7.804
Mean NS migration cost -7.943 -5.633 -2.609 -3.456
Mean NS delay -5.328 -4.082 -5.452 -4.349
Mean NS storage cost -4.994 -4.983 0.000 0.000
Mean NS computing cost -0.910 -1.043 -0.932 -1.003
Shared service profiles: In this scenario (Fig. 7), we consider multiple users sharing the same service profile (SP). We model 100 users, with 70 users using SP Type 1, and 30 using SP Type 2. These users are spread over a 2-city map, with 2APs in one city and 3 in another. It has a failure cost of 1000 and failure rate of 1e-2 [44]. If a user’s required AP is in another city, the delay cost between cities will be nearly as large as the failure cost, which is larger than the delay cost when accessing an SP within the same city. We apply FIRE-ImDQL to this scenario and compare its performance with NIS, RES and the greedy baseline. The greedy baseline deploys SPs to the $N$ most probable places w.r.t distribution of user transitions, where we search through all $N$ and plot the results with the lowest costs.
@@
Future work: Scalability to a large number of users: If users do not share service profiles, resource allocation is mostly de-coupled, except for the ‘queuing delay at server’ cost. Under this setting, the algorithm may not be as scalable for larger state spaces, because in our importance sampling algorithm, we update the functions $T(s)$ and $U(s)$ at state level, necessitating us to sample each state multiple times. Future areas of investigation can include modifying the importance sampling-RL algorithm (which is trained offline), to sample multiple trajectories in parallel while updating $T$ and $U$, for faster convergence. Another future area of investigation is to do “intelligent” or dynamic decoupling of the state space over base stations, via splitting the network into smaller subsets of groups of APs based on user mobility patterns. This makes the state space smaller. The algorithm can then be run on each ‘smaller state space’.